# On DR-semigroups satisfying the ample conditions

### v1

@@ 4.1 Background on ESN theorems
  • $es$ is the unique $t\leq_{r}s$ for which $D(t)=e$ (the “restriction of $s$ to $e$”), and
  • $sf$ is the unique $t\leq_{l}s$ for which $R(t)=f$ (the “co-restriction of $s$ to $f$”).
@@
$xy=x|e\circ e|y,\mbox{ where }e=R(x)D(y),$
meaning that the original semigroup product and hence the full structure of $S$ as an Ehresmann semigroup is captured by its derived biordered category $(S,\circ,D,R),\leq_{r},\leq_{l})$.
@@ 4.2 The (cat,trace)-product and partial categories
The resulting structure $(S,\circ,D,R)$ is evidently a category if and only if $S$ satisfies the cat-semigroup condition as in [10], which we may state as “$cat(x,y)\Rightarrow trace(x,y)$”.
@@
Let $\circ$ be the (cat,trace)-product on $S$. If $x\circ y$ exists then $R(x)=D(y)$ (but not necessarily conversely!), and then also $D(x\circ y)=D(x)$, and $R(x\circ y)=R(y)$, and the laws $D(x)\circ x=x\circ R(x)=x$ are easily seen to hold. However, $(S,\circ)$ may not be a partial semigroup. Stein was led to define the notion of a category with partial products in [7], as follows: it is a structure $C=(C,\circ,D,R)$, equipped with a partial binary operation $\circ$ and two unary operations $D$ and $R$ such that
  1. (PC1)
    $D(x)\circ x=x$, $x\circ R(x)=x$
  2. (PC2)
    $R(D(x))=D(x)$, $D(R(x))=R(x)$
  3. (PC3)
    if $x\circ y$ exists then $D(x\circ y)=D(x),R(x\circ y)=R(y)$ and $R(x)=D(y)$
  4. (PC4)
    $(C,\circ)$ is a partial semigroup.
The category case arises by further requiring that $x\circ y$ exists if and only if $R(x)=D(y)$. Note that Law (PC3) above no longer follows from the other laws as it did in the category case. Elements of $D(C)$ will be called identities as in the category case.
For ease of reference, we call categories with partial products partial categories in what follows. (Categories already have only a partial product, so we prefer our name, which is suggestive of the fact that $R(x)=D(y)$ is only necessary for $x\circ y$ to exist, not sufficient.)
As noted above, $(S\circ,D,R)$ satisfies all but perhaps Law (PC4), if $\circ$ is the (cat,trace)-product. In [7], Stein found a suitable sufficient condition on the DR-semigroup $S$, not necessarily satisfying the congruence conditions, to ensure that the derived structure $(S,\circ,D,R)$ is a partial category, or equivalently, to ensure that $(S,\circ)$ is a partial semigroup. His condition, which was that one of his generalized ample conditions hold (see Lemma 3.8 in [7]), was dictated by representation theory considerations: when is the semigroup algebra of a DR-semigroup isomorphic to the algebra obtained from its derived partial algebra under the (cat,trace)-product?
@@
Another problem is that $D(xy)x$ contains the term “$xy$”. If the congruence conditions are assumed, then we may write $D(xy)=D(x(R(x)D(y)))$, and then we can express $D(xy)$ in the language of Ehresmann categories, bur it is already known how to view Ehresmann (hence restriction) semigroups as Ehresmann categories; see [4].
Fortunately, we do have the following, mitigating the need to consider restrictions and corestrictions entirely.
@@
Suppose $S$ satisfies the ample conditions. Then for all $x,y\in S$, $D(xy)x$ is the largest $x_{1}\in S$ and $yR(xy)$ is the largest $y_{1}\in S$ (both under $\leq$) for which $x_{1}\leq x,y_{1}\leq y$ and $x_{1}\circ y_{1}$ exists under the (cat,trace)-product.
@@
Conversely, if $x_{1}\leq x,y_{1}\leq y$ and $x_{1}\circ y_{1}$ exists, we have $xy_{1}=xD(y_{1})y_{1}=xR(x_{1})y=x_{1}y_{1}$, and so
@@ 4.3 Ample partial categories
Suppose $x\leq z,y\leq z$ and $D(x)\leq D(y)$. Now there are largest $e\leq D(y),z^{\prime}\leq z$ such that $e\circ z^{\prime}$ exists (whence $e=D(z^{\prime})$ since $e\in D(C)$ by the first part). Since $D(y)\leq D(y),y\leq z$ and $D(y)\circ y$ exists, it must be that $D(y)\leq e\leq D(y)$, so $D(y)=e=D(z^{\prime})$, and $y\leq z^{\prime}$, and so by (APC6), $y=z^{\prime}$. Similarly there are largest $D(x)^{\prime}\leq D(x),z^{\prime\prime}\leq z$ such that $D(x)^{\prime}\circ z^{\prime\prime}$ exists. Arguing as above, $D(x)^{\prime}=D(x)$ and $x=z^{\prime\prime}$. Now $D(x)\leq D(y)$, $x\leq z$ and $D(x)\circ x$ exists, so by maximality of $D(y),y$ such that $D(y)\leq D(y)$, $y\leq z$ and $D(y)\circ y$ exists, we must have that $x\leq y$. Dually for the third property. $\Box$
@@
    For all $y\in C$ and $e\leq D(y)$, there is at most one $x\leq y$ with $D(x)=e$, the restriction of $x$ by $e$;
@@
    for all $y\in C$ and $e\leq R(y)$, there is at most one $x\leq y$ with $R(x)=e$, the corestriction of $x$ by $e$;
@@
Proof.    The first two claims are immediate from (APC6) and (APC7). For the third, picking $x^{\prime},y^{\prime}$ as in (APC5) and letting $e=R(x^{\prime})=D(y^{\prime})$, clearly $x^{\prime}=x|e,y^{\prime}=e|y$ and so $x|e\circ e|y$ exist. If also $x|f,f|y$ and $x|f\circ f|y$ exist then by maximality of $x^{\prime},y^{\prime}$ we have $x|f\leq x|e,f|y\leq e|y$, so $f=R(x|f)\leq R(x|e)=e$. $\Box$
@@ 4.4 The ESN-style theorem
$x\otimes y=x^{\prime}\circ y^{\prime},\mbox{ where $x^{\prime}\leq x,y^{\prime}\leq y$ are largest such that $x^{\prime}\circ y^{\prime}$ exists},$
@@
Conversely, suppose $(C,\circ,D,R,\leq)$ is an ample partial category. Define $s*t=s\otimes t$ whenever $cat(s,t)$ and $trace(s,t)$ holds in $(S,\otimes,D,R)$; this is the partial category product in $C(S(C))$. Now for all $s,t\in C$, $s*t$ exists in $C(S(C))$ if and only if $R(s)=D(t)$, $D(s\otimes t)=D(s)$ and $R(s\otimes t)=R(t)$. So if $s*t$ exists then writing $s\otimes t=s^{\prime}\circ t^{\prime}$ for some $s^{\prime}\leq s,t^{\prime}\leq t$ in $C$, we have that $D(s^{\prime}\circ t^{\prime})=D(s)$, so $D(s^{\prime})=D(s)$ and so $s^{\prime}=s$ by (APC6) for ample partial categories, and dually $t^{\prime}=t$, so $s*t=s\otimes t=s\circ t$. Conversely, if $s\circ t$ exists then obviously $s\otimes t=s\circ t$, so $R(s)=D(t)$, $D(s\otimes t)=D(s)$ and $R(s\otimes t)=R(t)$, and so $s*t$ exists. Again, $D$ and $R$ are unchanged in passing from $C$ to $C(S(C))$.
@@
Proof.    Suppose $S,T$ are DR-semigroups satisfying the ample conditions and that $f:S\rightarrow T$ is a homomorphism respecting $D,R$. If $a,b\in S$ are such that $a\circ b$ exists in $C(S)$, then $cat(a,b)$ and $trace(a,b)$ hold, so applying the homomorphism property of $f$ we obtain $cat(f(a),f(b))$ and $trace(f(a),f(b))$ hold, and so $f(a)\circ f(b)$ exists and evidently equals $f(a)f(b)=f(ab)=f(a\circ b)$, again from the homomorphism property of $f$. Of course, $f$ respects $D$ and $R$. Moreover, if $a,b\in S$ are such that $a\leq b$ then $D(a)\leq D(b)$, so $D(a)=D(a)D(b)$ and so $D(f(a))=f(D(a))=f(D(a)D(b))=D(f(a))D(f(b))$, so $D(f(a))\leq D(f(b))$, but also $a=D(a)b$, so $f(a)=D(f(a))f(b)$, so $f(a)\leq f(b)$. So $f$ can be viewed as a functor $C(S)\rightarrow C(T)$ that respects $\leq$. Furthermore, we know that $x^{\prime}=D(xy)x=xR(D(xy)x)$ and $y^{\prime}=yR(xy)=D(yR(xy))y$ are the largest $x^{\prime}\leq x,y^{\prime}\leq y$ for which $x^{\prime}\circ y^{\prime}$ exists by Proposition 4.4, and that the matching term for $x,y$ is $e=R(D(xy)x)=D(yR(xy))$; similarly, the matching term for $f(x),f(y)$ in $C(T)$ is
$R(D(f(x)f(y))f(x))=R(D(f(xy))f(x))=R(f(D(xy)x))=f(R(D(xy)x))=f(e).$
@@
Conversely, suppose $g:C_{1}\rightarrow C_{2}$ is an ample functor. Of course, $D$ and $R$ are preserved, and for all $s,t\in S$, $s\otimes t=s^{\prime}\circ t^{\prime}$, where $s^{\prime}\leq s,t^{\prime}\leq t$ are largest such that $s^{\prime}\circ t^{\prime}$ exists. Then $g(s\otimes t)=g(s^{\prime}\circ t^{\prime})=g(s^{\prime})\circ g(t^{\prime})=g(s)\otimes g(t)$ by the previous proposition and the definition of $\otimes$. This shows that we have a functor $G$ from the category of ample partial categories to the category of DR-semigroups satisfying the ample conditions in which $G(C)=S(C)$ and for the ample functor $g:C_{1}\rightarrow C_{2}$, $G(g):S(C_{1})\rightarrow S(C_{2})$ is defined by setting $G(g)=g$ as functions.
@@
As seen in Section 3.2, the DR-semigroup of binary relations under double demonic composition satisfies the ample and cat-semigroup conditions, so the derived ample partial category is in fact a category. For this case, the only change needed is to replace “partial category” by “category” in the definition of ample partial categories, thereby obtaining the class of ample categories. The above theorem then specialises to give an isomorphism between the category of DR-semigroups satisfying the cat-semigroup and ample conditions on the one hand and the category of ample categories on the other.
@@ 5 An embedding theorem
Let $C$ be a partial category such that $D(C)$ is partially ordered, and let $P(C)$ consist of all subsets of $C$. (Evidently, $C$ could be a category in which the partial order on $D(C)$ is equality.) Because $C$ is a partial semigroup, $P(C)$ can obviously be made into a semigroup by defining, for all $A,B\in P(C)$,
$AB=\{a\circ b\mid a\in A,b\in B\}.$
It is well-known that if $C$ is a category and the partial order on $D(C)$ is taken to be equality, then $(P(C),\cdot,D,R)$ is an Ehresmann semigroup, generalizing the fact that the set of binary relations on a set gives an Ehresmann semigroup under relational composition, domain and range. More generally, we have the following.
Proposition 5.1
Proof.    Certainly $(P(C),\cdot)$ is a semigroup.
Proposition 5.2
Proof.    Suppose $A\in B(C)$. Pick $a,b\in A$. Suppose $D(a)\leq D(b)$. Now let $E=\{e\in D(C)\mid e\leq R(b)\}$. Then $E=D(E)$, and so because $A\in B(C)$, $AE=D(AE)A$. Hence, as $b=b\circ R(b)\in AE$, we have $b\in D(AE)A$, so $b=e\circ a$ for some $e\in D(AE),a\in A$, meaning that $e=D(a),b=a$, so $a\in D(AE)A=AE$, and so $a=a^{\prime}\circ e$ for some $e\in E$, and so $a=a^{\prime},e=R(a)\in E$ and so by definition of $E$, $R(a)\leq R(b)$. Dualising gives the converse.
Proposition 5.3

### v2

@@ 4.1 Background on ESN theorems
  • eses is the unique $t\leq_{r}s$ for which $D(t)=e$ (the “restriction of $s$ to ee”), and
  • sfsf is the unique $t\leq_{l}s$ for which $R(t)=f$ (the “co-restriction of $s$ to ff”).
@@
$xy=(xe)\circ(ey),\mbox{ where }e=R(x)D(y),$
meaning that the original semigroup product and hence the full Ehresmann semigroup structure of $S$ is captured by its derived biordered category $(S,\circ,D,R,\leq_{r},\leq_{l})$.
@@ 4.2 The (cat,trace)-product and partial categories
The resulting structure $(S,\circ,D,R)$ is evidently a category if and only if $S$ satisfies the cat-semigroup condition as in [11], which we may state as $cat(x,y)\Rightarrow trace(x,y)$”.
@@
Let $\circ$ be the (cat,trace)-product on $S$. If $x\circ y$ exists then $R(x)=D(y)$ (but not necessarily conversely!), and then also $D(x\circ y)=D(x)$, and $R(x\circ y)=R(y)$, and the laws $D(x)\circ x=x\circ R(x)=x$ are easily seen to hold. However, $(S,\circ)$ may not be a partial semigroup.
In [8], Stein found a suitable sufficient condition on the DR-semigroup $S$, not necessarily satisfying the congruence conditions, to ensure that the derived structure $(S,\circ,D,R)$ is a partial category, or equivalently, to ensure that $(S,\circ)$ is a partial semigroup. His condition, which was that one of his generalized ample conditions hold (see Lemma 3.8 in [8]), was dictated by representation theory considerations: when is the semigroup algebra of a DR-semigroup isomorphic to the algebra obtained from its derived partial algebra under the (cat,trace)-product?
@@
Another problem is that $D(xy)x$ contains the term “xyxy”. If the congruence conditions are assumed, then we may write $D(xy)=D(x(R(x)D(y)))$, and then we can express $D(xy)$ in the language of Ehresmann categories, but it is already known how to view Ehresmann (hence restriction) semigroups as Ehresmann categories; see [5].
Fortunately, we do have the following, mitigating the need to consider restrictions and corestrictions entirely. First, recall the definition of the standard order on the DR-semigroup $S$ satisfying the ample conditions, as in the second part of Proposition 2.1.
@@
Suppose $S$ satisfies the ample conditions. Then for all $x,y\in S$, $D(xy)x$ is the largest $x_{1}\in S$ and $yR(xy)$ is the largest $y_{1}\in S$ (both under $\leq$) for which $x_{1}\leq x$ and $y_{1}\leq y$ and for which $x_{1}\circ y_{1}$ exists under the (cat,trace)-product.
@@
Conversely, if $x_{1}\leq x,y_{1}\leq y$ and $x_{1}\circ y_{1}$ exists, we have $xy_{1}=xD(y_{1})y_{1}=xR(x_{1})y_{1}=x_{1}y_{1}$, and so
@@ 4.3 Ample partial categories
Suppose $x\leq z,y\leq z$ and $D(x)\leq D(y)$. Now there are largest $e\leq D(y),z^{\prime}\leq z$ such that $e\circ z^{\prime}$ exists (whence $e=D(z^{\prime})$ since $e\in D(C)$ by the first part). Since $D(y)\leq D(y),y\leq z$ and $D(y)\circ y$ exists, it must be that $D(y)\leq e\leq D(y)$, so $D(y)=e=D(z^{\prime})$, and $y\leq z^{\prime}$, and so by (APC6), $y=z^{\prime}$. Now $D(x)\leq D(y)$, $x\leq z$ and $D(x)\circ x$ exists, so by maximality of $D(y),y$ such that $D(y)\leq D(y)$, $y\leq z$ and $D(y)\circ y$ exists, we must have that $x\leq y$. Dually for the third property. $\Box$
@@
The parenthetical uniqueness claim made in Law (APC8) above now follows, for if $x^{\prime}\circ y^{\prime}=x^{\prime\prime}\circ y^{\prime\prime}$ where $x^{\prime},x^{\prime\prime}\leq x$ and $y^{\prime},y^{\prime\prime}\leq y$, then $D(x^{\prime})=D(x^{\prime}\circ y^{\prime})=D(x^{\prime\prime}\circ y^{\prime\prime})=D(x^{\prime\prime})$ and so by the previous corollary, $x^{\prime}=x^{\prime\prime}$, and dually $y^{\prime}=y^{\prime\prime}$.
@@
    For all $y\in C$ and $e\leq D(y)$, there is at most one $x\leq y$ with $D(x)=e$, the restriction of $y$ by $e$, denoted by $e|y$;
@@
    for all $y\in C$ and $e\leq R(y)$, there is at most one $x\leq y$ with $R(x)=e$, the corestriction of $y$ by $e$, denoted by $y|e$;
@@
Proof.   The first two claims are immediate from Corollary 4.6. For the third, picking $x^{\prime},y^{\prime}$ as in (APC5) and letting $e=R(x^{\prime})=D(y^{\prime})$, clearly $x^{\prime}=x|e,y^{\prime}=e|y$ and so $x|e\circ e|y$ exist. If also $x|f,f|y$ and $x|f\circ f|y$ exist then by maximality of $x^{\prime},y^{\prime}$ we have $x|f\leq x|e,f|y\leq e|y$, so $f=R(x|f)\leq R(x|e)=e$. $\Box$
@@ 4.4 The ESN-style theorem
[…]
@@
Conversely, suppose $(C,\circ,D,R,\leq)$ is an ample partial category. Define $s*t=s\otimes t$ whenever $cat(s,t)$ and $trace(s,t)$ holds in $(S,\otimes,D,R)$; this is the partial category product in $C(S(C))$. Now for all $\in C, sts*t$ exists in $C(S(C))$ if and only if $R(s)=D(t)$, $D(s\otimes t)=D(s)$ and $R(s\otimes t)=R(t)$. So if sts*t exists then writing $s\otimes t=s^{\prime}\circ t^{\prime}$ for some $s^{\prime}\leq s,t^{\prime}\leq t$ in $C$, we have that $D(s^{\prime}\circ t^{\prime})=D(s)$, so $D(s^{\prime})=D(s)$ and so $s^{\prime}=s$ by (APC6) for ample partial categories, and dually $t^{\prime}=t$, so $s*t=s\otimes t=s\circ t$. Conversely, if $s\circ t$ exists then obviously $s\otimes t=s\circ t$, so $R(s)=D(t)$, $D(s\otimes t)=D(s)$ and $R(s\otimes t)=R(t)$, and so sts*t exists. Again, $D$ and $R$ are unchanged in passing from $C$ to $C(S(C))$.
@@
The following is a generalisation of Theorem 5.7 in [5], where restriction semigroups were called idempotent-connected semigroups and the relevant class of categories was called the class of inductive1-categories.
@@
Proof.   Suppose $S,T$ are DR-semigroups satisfying the ample conditions and that $f:S\rightarrow T$ is a homomorphism respecting $D,R$. If $a,b\in S$ are such that $a\circ b$ exists in $C(S)$, then $cat(a,b)$ and $trace(a,b)$ hold, so applying the homomorphism property of $f$ we obtain $cat(f(a),f(b))$ and $trace(f(a),f(b))$ hold, and so $f(a)\circ f(b)$ exists and evidently equals $f(a)f(b)=f(ab)=f(a\circ b)$, again from the homomorphism property of $f$. Of course, $f$ respects $D$ and $R$. Moreover, if $a,b\in S$ are such that $a\leq b$ then $D(a)\leq D(b)$, so $D(a)=D(a)D(b)$ and so D(f(a))=f(D(a))=f(D(a)D(b))=D(f(a))D(f(b))D(f(a))=f(D(a))=f(D(a)D(b))=D(f(a))D(f(b)), so $D(f(a))\leq D(f(b))$, but also $a=D(a)b$, so $f(a)=D(f(a))f(b)$, so $f(a)\leq f(b)$. So $f$ can be viewed as a functor $C(S)\rightarrow C(T)$ that respects $\leq$. Furthermore, we know that $x^{\prime}=D(xy)x=xR(D(xy)x)$ and $y^{\prime}=yR(xy)=D(yR(xy))y$ are the largest $x^{\prime}\leq x,y^{\prime}\leq y$ for which $x^{\prime}\circ y^{\prime}$ exists by Proposition 4.4, and that the matching term for $x,y$ is $e=R(D(xy)x)=D(yR(xy))$; similarly, the matching term for $f(x),f(y)$ in $C(T)$ is
R(D(f(x)f(y))f(x))=R(D(f(xy))f(x))=R(f(D(xy)x))=f(R(D(xy)x))=f(e).R(D(f(x)f(y))f(x))=R(D(f(xy))f(x))=R(f(D(xy)x))=f(R(D(xy)x))=f(e).
@@
[…]
@@ 5 An embedding theorem
Acknowledgements
The author is very grateful to the anonymous referee for comments that led to significant improvements in the content and layout of the article, especially the first and third sections.